# Supervised Fine-Tuning: An Activation Pattern Optimization Process for Attention Heads

### v1

@@ Abstract
Though demonstrating promising potential, LLMs’ performance on complex tasks, such as advanced mathematics and complex disease diagnosis is still unsatisfactory. A key issue is the present LLMs learn in a data-driven schema, while the instruction dataset about these complex tasks is both scarce and hard to collect or construct. On the contrary, a prominent phenomenon is that LLMs can learn rather fast on those simpler tasks with adequate prior knowledge captured during pretraining stage. Thus, if the prerequisite and mechanism of such rapid generalization could be elucidated, it could be highly beneficial in enhancing the efficiency and effectiveness of the LLM’s ability to learn complex tasks. Thus, in this paper, we employ a gradient-based method, to dissect the process that the SFT process adapts LLMs to downstream tasks via the perspective of attention patterns. We find that: (1) LLMs selectively activate task-specific attention heads during SFT; (2) activation patterns for complex tasks are combinations of basic task patterns; and (3) changes in a few parameters can significantly impact activation patterns after SFT on a small number of samples. Based on these insights, we conduct experiments to examine whether these conclusions could effectively enhance the efficiency and effectiveness of SFT, particularly in handling complex tasks and when instructional resources are scarce. Our research not only uncovers the underlying reasons behind LLMs’ rapid learning and generalization mechanisms but also provides practical solutions for addressing data challenges in complex and specialized tasks.
Introduction
The Supervised Fine-tuning Process (SFT) is essential for aligning Large Language Models (LLMs) with human objectives to complete downstream tasks (Zhang et al. 2023). Leveraging the extensive prior knowledge acquired during pretraining, fine-tuning LLMs on instruction sets enables them to perform well across various tasks. (Dong et al. 2023; Xia et al. 2024). However, in more complex tasks like mathematical reasoning, LLM performance remains unsatisfactory, indicating there is still considerable room for improvement. (Fourrier et al. 2024; Contributors 2023).
One major reason for such performance limitation is that these complex tasks often demand simultaneous utilization of multiple types of knowledge and skills, and thus, more instructions are required to adapt LLMs to these tasks (Kaplan et al. 2020). However, the more complex the task, the more challenging it is to collect and construct relevant instruction data (Zhang et al. 2023; Minaee et al. 2024). This in turn limits the ability of LLMs to improve performance on these tasks. For instance, data related to complex decision-making processes, such as medical decision-making and social behavior prediction, are often challenging to construct or obtain due to their sensitive nature, proprietary status, or the prohibitively high costs associated with data collection.
@@ Introduction
In contrast, on most basic tasks, LLMs demonstrate prominent rapid generalization (Zhang et al. 2023). By training on just a few thousand instructions, LLMs can learn to complete various tasks(Xia et al. 2024). Therefore, understanding the mechanisms and conditions that enable LLMs’ rapid learning and generalization could strongly guide their adaptation to complex tasks.
@@
To address this issue, we propose to analyze how SFT adapts LLMs to different tasks through the perspective of activation patterns of attention heads using a gradient-based method. Previous studies show that attention heads serve as basic functional units in transformer-based models (Voita et al. 2019; Clark et al. 2019; Hao et al. 2021; Wang et al. 2023). Through pretraining, these attention heads capture different types of information, model various relationships, and demonstrate distinct functions across tasks.
@@
Therefore, analyzing how LLMs utilize different functions to solve tasks during the SFT process could elucidate the mechanisms behind SFT learning and generalization.If an attention head influences outputs in certain tasks, it is heuristically considered “activated”. The composition of activations across different attention heads forms the activation pattern of the LLMs (Figure 1), indicating how the model’s functional units are integrated to solve a specific task.To determine if an attention head influences the output (i.e., is activated), we use a gradient-based method because gradients naturally measure the sensitivity and impact of outputs on input features  (Wang et al. 2024). By calculating the expected value of the gradient on the attention score matrix, we can quantify the impact of each attention head on model performance before and after SFT This approach reveals the changes in LLMs following SFT, and explains how SFT adapts LLMs to downstream tasks.
@@
By analyzing the alternation of attention head activation patterns upon tasks across multiple domains during the SFT process, we find the critical role of the change of activation pattern in fast learning and generalization during the SFT process, as: (1) LLMs learns to complete tasks by selectively activate attention heads during the SFT process. After SFT, the LLMs activates certain tasks-specific attention heads, and the differences in activation patterns between tasks become more pronounced; (2) LLMs complete tasks by learning to integrate basic skills. We observe that the activation patterns on completing complex tasks are a combination of several basic tasks; (3) Changes in a few parameters can significantly impact activation patterns after SFT on a small number of samples. The model’s activation patterns exhibit significant changes compared to the base model in the early stage of the SFT process with rather limited optimization steps, showing that minimal parameter adjustments can significantly alter activation patterns.
@@
Based on these findings, we further validate them by examining whether they can enhance the effectiveness and efficiency of SFT, especially on complex tasks and when instructions are scarce. We conduct two experiments: The first simulates a scenario with limited complex task data. By fine-tuning LLMs using instructions on the basic skills required for these tasks, LLM performance significantly improves. The second experiment addresses scenarios where high-quality domain data is private or unavailable. Based on activation patterns, we select relevant instructions from a large pool to approximate the effects of private data using public instructions, thereby improving LLM performance. These results not only support our findings on rapid learning and generalization mechanisms, but also offer practical solutions for overcoming data challenges in complex and specialized tasks.
Background and Related Work
One factor limiting the effectiveness of LLMs is the opacity of their mechanisms and the lack of understanding of generalization in SFT. Some studies have used gradient-based methods to identify key features and parameters influencing LLM outputs, shedding light on their internal mechanisms (Hao et al. 2021; Wang et al. 2023; Yang, Wang, and Bilgic 2023). However, most research on SFT treats LLMs as black boxes (He et al. 2024; Dong et al. 2023; Lu et al. 2023). Given this black-box approach, there is a gap in understanding the detailed mechanisms behind SFT, which our work aims to address. Our work takes a pioneering approach in this area by exploring these underlying mechanisms.
A key question is which perspective is most effective for this analysis. Previous studies (Yu et al. 2024; Fu et al. 2023) suggest that parameter changes after SFT are limited, and given the vast number of parameters, interpreting their impact is challenging.
@@
In contrast, previous studies have identified that an attention head can act as a basic functional unit in transformer-based models (Voita et al. 2019; Clark et al. 2019; Wang et al. 2023), with different attention heads serving distinct functions. For example, Voita et al. (2019) and Clark et al. (2019) found that attention heads at different layers focus on different parts-of-speech, while Wang et al. (2023) observed that during in-context learning, attention heads at different layers focus on different parts of the in-context learning process. Additionally, prior research (Jin et al. 2024; Lee and Hwang 2024) has suggested that attention heads exhibit task specificity, as pruning certain attention heads does not lead to performance degradation on downstream tasks. Heuristically, since a task can be viewed as a combination of several basic tasks, completing it requires the coordination of multiple corresponding basic functional units, i.e., attention heads. Thus, in this paper, we investigate the fast generalization mechanism of SFT from the perspective of activation patterns in attention heads, using a gradient-based method for our analyses..
@@
For clarity, we define attention heads that impact downstream tasks as activated and those that do not as inactive. The matrix representing the influence of attention heads on downstream tasks is called the activation pattern. By examining how attention head activations change in response to fine-tuning, we aim to uncover the mechanisms enabling rapid learning and generalization in LLMs.
Methodology
@@
During pretraining, LLMs learn to use different attention heads to capture various types of information (Voita et al. 2019; Khaki and Plataniotis 2024). Heuristically, a complex task can often be viewed as a composition of several basic tasks. For instance, solving a mathematical problem using code, as in SGSM (Christ, Kropko, and Hartvigsen 2024), can be broken down into fundamental tasks like code generation and mathematical reasoning. Since each attention head may correspond to a specific function, an LLM can invoke multiple attention heads and integrate their functions to accomplish a complex task. Combining these observations leads to our core assumption: during the SFT process, LLMs quickly adapt to downstream tasks by invoking different attention heads. Thus, in this paper, we investigate the mechanism of fast generalization in the SFT process by analyzing changes in the activation patterns of attention heads.
Gradient-Based Attention Head Activation Patterns Analysis
@@
One critical issue is how to measure if an attention head has been used in a task, in other words, it has been activated. Inspired by (Hao et al. 2021) and (Wang et al. 2023), we measure this using model gradients, as gradients can naturally reflect the subtle effects of input features or parameters on the model’s output (Wang et al. 2024). When the gradient of a particular attention head is large, it indicates that the model’s output is more sensitive to the attention head, suggesting a higher influence on outputs.
@@
Specifically, given an LLMs with $L$ layers and $H$ attention heads per layer, on a dataset $\mathcal{T}$ about one certain task, the activation pattern (AP) could be represented as a $L\times H$ matrix, with each element representing the activation level (AL) reflecting the contribution of one specific attention head to a given task:
@@
where $N$ represents size $\mathcal{T}$, $x_{i}$ is the $i$th instance in $\mathcal{T}$ and $L(x_{i})$ is the loss value of the model on $x_{i}$. $\Gamma_{l,h}$ is the attention matrix of the $hl$th attention head. $\partial\Gamma_{l,h}$ is the gradient of the $lh$th attention head with respect to $L(x_{i})$. By combining the absolute value of attention scores with the sensitivity of loss on the attention scores, the total influence of a certain attention head on outputs and the task could be quantified. Thus, $AP^{task}$, composed of $AL^{task}$, reflects how each attention head contributes to task completion.
Analytical Framework
@@
To validate our core assumption, we focus on three progressive issues: (1) How do the activation patterns change during the SFT process for each task, and are these changes specific to the task? In other words, can we grasp the characteristics of the task from the perspective of activation patterns? (2) What is the relationship between the activation patterns when learning complex tasks and those when learning basic tasks? This reflects the prerequisites that may be necessary for learning complex tasks; and (3) on basic tasks, how many training samples are required to significantly change the activation pattern? The necessary sample size indicates the model’s ability to achieve rapid generalization.
@@
First, we evaluated changes in activation patterns before and after SFT across tasks. We observed that more attention heads were activated post-SFT, with these activations being task-specific, suggesting that during the SFT Process, Attention Heads are Selectively Activated. Next, we explored the relationship between activation patterns in complex and basic tasks, finding that the patterns of several basic tasks can effectively approximate those of complex tasks. From this, we concluded that activation patterns in complex tasks are combinations of basic task patterns. Finally, by dynamically observing activation pattern changes during SFT, we concluded that even small datasets can significantly alter activation patterns.
@@
To further validate our conclusions, we applied our findings to enhance the effectiveness of SFT on complex tasks. By evaluating the resulting improvements, we further confirmed the validity of our theory.The following sections provide a detailed analysis of the process and the conclusions drawn from the analysis.
SFT Adapts to Downstream Tasks by Modifying Activation Patterns
This section employs gradient-based analysis to examine the impact of SFT on transformer attention head activation patterns. Our analysis produces three key conclusions that address the questions previously posed.
Firstly, after the SFT process, LLMs learn to complete tasks by selectively activating attention heads. Secondly, changes in activation patterns for compound tasks can be interpreted as a combination of changes observed in more basic tasks.Thirdly, during the SFT process, the model’s activation patterns exhibit exponential changes, meaning that even a small amount of data can significantly alter these patterns.
Attention Heads are Selectively Activated
@@
In this section, we first analyze the changes in attention head activation patterns before and after SFT across different tasks. Then, we explore the relationship between these activation pattern changes across tasks.
Analytical Method
@@
To analyze the changes in model activation patterns before and after SFT for each task, we used three metrics: Gini Coefficient, Coefficient of Variation (CV), and Kurtosis. The Gini Coefficient measures distribution inequality, with values above 0.4 indicating significant concentration in a few attention heads. The Coefficient of Variation reflects relative dispersion, with values over 1 indicating substantial variability. Kurtosis describes the distribution’s shape, where values above 3 suggest heavy tails and sharp peaks, indicating the presence of extreme values.
To analyze the relationship between changes in activation patterns across different tasks, we used the Pearson correlation coefficient of activation changes, which reflects the linear consistency of the direction of changes. A high correlation between the activation patterns of two tasks indicates that the attention heads relied upon to complete both tasks are highly similar. Additionally, a high correlation in the rate of change of activation patterns suggests that the degree to which attention heads are engaged during the SFT process is similar for both tasks.
Experiments Setup
@@
We conducted experiments on models with varying performance levels, including Llama3-8B (AI@Meta 2024), Gemma-7B (Team et al. 2024), and OPT-6.7B (Zhang et al. 2022), across a range of tasks: mathematical reasoning (MATH (Hendrycks et al. 2021b), GSM8K (Cobbe et al. 2021)), coding (Code Search Net (Husain et al. 2019), SGSM (Christ, Kropko, and Hartvigsen 2024)), and natural language processing and reasoning (HellaSwag (Zellers et al. 2019), Winogrande (Sakaguchi et al. 2021), ARC (Clark et al. 2018)). Details on the data and training process are provided in Appendix I and II.
Results

### v2

# Supervised Fine-Tuning Achieve Rapid Task Adaption Via Alternating Attention Head Activation Patterns

@@ Abstract
LLMs’ performance on complex tasks is still unsatisfactory. A key issue is that presently LLMs learn in a data-driven schema, while the instructions about these complex tasks are both scarce and hard to collect or construct. On the contrary, a prominent phenomenon is that LLMs can learn rather fast on simpler tasks with adequate prior knowledge captured during pretraining stage. Thus, if the prerequisite and mechanism of such rapid generalization could be elucidated, it could enhance the efficiency and effectiveness of the LLM’s ability to learn complex tasks. Thus, in this paper, we employ a gradient-based method, to dissect the process that the SFT process adapts LLMs to downstream tasks via the perspective of attention patterns. We find that: (1) LLMs selectively activate task-specific attention heads during SFT; (2) activation patterns for complex tasks are combinations of basic task patterns; and (3) changes in a few parameters can significantly impact activation patterns after SFT on a small number of samples. Based on these insights, experiments are conducted to actually enhance the efficiency and effectiveness of SFT.
Supervised Fine-Tuning Achieve Rapid Task Adaption Via Alternating Attention Head Activation Patterns
Yang Zhao [^1] These authors contributed equally to this work., Li Du  [^1] These authors contributed equally to this work.,  Xiao Ding [^2] Corresponding Author.,   Kai Xiong,   Ting Liu and Bing Qin Research Center for Social Computing and Information Retrieval Harbin Institute of Technology, China Beijing Academy of Artificial Intelligence, Beijing, China {yangzhao, xding, kxiong, tliu, qinb}@ir.hit.edu.cn duli@baai.ac.cn ,
1 Introduction
The Supervised Fine-tuning Process (SFT) is essential for optimizing Large Language Models (LLMs) to effectively complete downstream tasks  Zhang et al. (2023). Leveraging the extensive prior knowledge acquired during pretraining, fine-tuning LLMs on instruction sets enables them to perform well across various tasks.  Dong et al. (2023); Xia et al. (2024). However, in more complex tasks like mathematical reasoning, LLM performance remains unsatisfactory.  Fourrier et al. (2024); Contributors (2023). One major reason for such performance limitation is that these complex tasks often require simultaneously using multiple types of knowledge and skills, and thus, more instructions are required to adapt LLMs to these tasks  Kaplan et al. (2020). However, the more complex the task, the more challenging it is to collect and construct relevant instruction data  Zhang et al. (2023); Minaee et al. (2024). This in turn limits the ability of LLMs to improve performance on these tasks.
@@ 1 Introduction
In contrast, on most basic tasks, LLMs demonstrate prominent rapid generalization Zhang et al. (2023). By training on just a few thousand instructions, LLMs can learn to complete various tasksXia et al. (2024). Therefore, understanding the mechanisms and conditions that enable LLMs’ rapid learning and generalization could strongly guide their adaptation to complex tasks.
@@
To address this issue, we propose to analyze the prerequisite and mechanism of such rapid task adaption during SFT from the perspective of activation patterns of attention heads using a gradient-based method. Previous studies show that attention heads serve as basic functional units in transformer-based models  Voita et al. (2019); Clark et al. (2019); Hao et al. (2021); Wang et al. (2023). These attention heads could capture different types of information and model various relationships for completing different tasks. Therefore, analyzing how LLMs learn to utilize these basic function units to solve tasks during the SFT process could shed light on the mechanisms behind instruction learning and generalization. If an attention head influences outputs in certain tasks, it is heuristically considered “activated”. The composition of activations across different attention heads forms the activation pattern of the LLMs (Figure 1), indicating how the model’s functional units are integrated to solve a specific task.
@@
To determine if an attention head influences the output, we use a gradient-based method because gradients naturally measure the sensitivity and impact of outputs on input features  Wang et al. (2024a). By calculating the expected value of the gradient on the attention score matrix, we can quantify the impact of each attention head on model performance before and after SFT and explains how SFT adapts LLMs to downstream tasks.
@@
From the perspective of attention head activation patterns, we find the critical role of the change of activation pattern in fast learning and generalization during the SFT process: (1) LLMs learn to complete tasks by selectively activating attention heads during the SFT process. After SFT, the LLMs activate certain tasks-specific attention heads, and the differences in activation patterns between tasks become more pronounced; (2) LLMs complete tasks by learning to integrate basic skills.This finding suggests that complex task-solving in LLMs may be decomposed into a series of simpler subtasks, offering a modular perspective on how LLMs can be more effectively fine-tuned for intricate problems; (3) Changes in a few parameters can significantly impact activation patterns after SFT on a small number of samples. The model’s activation patterns exhibit significant changes compared to the base model in the early stage of the SFT process with rather limited optimization steps, showing that minimal parameter adjustments can significantly alter activation patterns.
@@
We validate our findings by exploring their practical application in enhancing the effectiveness and efficiency of SFT, particularly for complex tasks and in scenarios where high-quality instructions are scarce. Specifically: (1) When complex task data is limited, by fine-tuning LLMs using instructions on the basic skills required for these tasks, the efficiency of instruction tuning significantly improves. (2) When high-quality domain data is private or unavailable, based on activation patterns, we can select relevant instructions from a large pool of publicly available data to approximate the effects of private data. These findings not only validate our understanding of rapid learning and generalization mechanisms but also provide a scalable framework for improving LLM performance in real-world, data-constrained scenarios, such as specialized industries or domains where data availability is limited.
2 Background and Related Work
One major factor limiting the learning efficiency of instruction learning of LLMs is the opacity of their internal mechanisms. Though pioneer studies recognized key features or parameters of LLM  Hao et al. (2021); Wang et al. (2023); Yang et al. (2023b), the mechanism of adapting LLMs towards downstream tasks during SFT is still largely unknown  He et al. (2024); Dong et al. (2023). In this paper, we aim to serve as a pioneer to explore these underlying mechanisms.
A critical issue is from which perspective should we start. Instead of focusing on the parameters of LLMs, we choose to investigate the activation pattern of attention heads, as:
@@ 2 Background and Related Work
(1) Previous studies have identified that an attention head can act as a basic functional unit in transformer-based models  Voita et al. (2019); Clark et al. (2019); Wang et al. (2023), with different attention heads serving distinct functions. For example,  Voita et al. (2019) and  Clark et al. (2019) found that attention heads at different layers focus on different parts-of-speeches, while  Wang et al. (2023) observed that during in-context learning, different attention heads focus on different parts of the prompts.
(2) Additionally, prior research  Jin et al. (2024); Lee and Hwang (2024) has suggested the task-specific activation of attention heads, as pruning certain attention heads would lead to task-specific performance degradation. Heuristically, since a task can be viewed as a combination of several basic tasks, completing it requires the coordination of multiple corresponding basic functional units, i.e., attention heads.
(3) In contrast, Previous studies Yu et al. (2024); Fu et al. (2023) have suggested that parameter changes after SFT are limited. Furthermore, given the vast number of parameters in LLMs, interpreting the significance of these parameters and their changes are highly challenging.
@@
For clarity, we define attention heads that impact downstream tasks as activated and those that do not as inactive. The matrix representing the influence of attention heads on downstream tasks is called the activation pattern.
3 Methodology
@@
During pretraining, LLMs learn to use different attention heads to capture various types of information  Voita et al. (2019); Khaki and Plataniotis (2024). Heuristically, a complex task can often be viewed as a composition of several basic tasks. For instance, solving a mathematical problem using code, can be broken down into fundamental tasks like code generation and mathematical reasoning. Since each attention head may correspond to a specific function, an LLM can invoke multiple attention heads and integrate their functions to accomplish a complex task. Combining these observations leads to our core assumption: during the SFT process, LLMs quickly adapt to downstream tasks by invoking different attention heads. Thus, in this paper, we investigate the mechanism of fast generalization in the SFT process by analyzing changes in the activation patterns of attention heads.
3.1 Gradient-Based Analysis of Attention Head Activation Patterns
@@
A key issue is determining whether an attention head is activated during a task. Inspired by Hao et al. (2021) and Wang et al. (2023), we measure this using model gradients, as they reflect the influence of input features or parameters on the model’s output  Wang et al. (2024a). A larger gradient for a particular attention head suggests that the model’s output is more sensitive to it, indicating a stronger influence on outputs.
@@
Specifically, given an LLM with $L$ layers and $H$ attention heads per layer, and a dataset $\mathcal{T}$, the activation pattern $AP^{\mathcal{T}}$ can be represented as an $L\times H$ matrix, with each element representing the activation level reflecting the contribution of one specific attention head to a given task:
@@
where $N$ represents the size of $\mathcal{T}$, $x_{i}$ is the $i$th instance in $\mathcal{T}$ and $L(x_{i})$ is the loss value of the model for $x_{i}$. $\Gamma_{l,h}$ is the attention matrix of the $h$ th attention head in the $l$ th layer.$\frac{\partial L(x_{i})}{\partial\Gamma_{l,h}}$ is the gradient of the $h$ th attention head in the $l$th layer with respect to $L(x_{i})$. By combining the absolute values of attention scores with the sensitivity of the loss to these scores, the total influence of a specific attention head on outputs and the task can be quantified. Thus, $AP^{task}$, composed of $AL^{task}$, reflects how each attention head contributes to task completion.
3.2 Analytical Framework
@@
To validate our core assumption, we focus on three progressive issues: (1) How do activation patterns change during the SFT process for each task, and are these changes task-specific? In other words, can we grasp the characteristics of the task from the perspective of activation patterns? (2) What is the relationship between activation patterns in complex tasks and basic tasks? This reflects the prerequisites that may be necessary for learning complex tasks; and (3) on basic tasks, how many training samples are required to significantly change the activation patterns? The necessary sample size indicates the model’s ability to achieve rapid generalization.
@@
First, we evaluated the changes in activation patterns across tasks before and after SFT. We observed that more attention heads are activated after SFT, and these activations are task-specific, suggesting that attention heads are selectively activated during SFT. Next, we explored the relationship between activation patterns in complex and basic tasks, finding that the patterns of several basic tasks effectively approximate those of complex tasks. This led us to conclude that activation patterns in complex tasks are combinations of basic task patterns. Finally, dynamic observations during SFT showed that even small datasets can significantly alter activation patterns.
@@
To further validate our conclusions, we applied these insights to enhance SFT effectiveness on complex tasks. The resulting improvements further confirmed our theory. The following sections provide a detailed analysis and conclusions.
4 SFT Adapts to Downstream Tasks by Modifying Activation Patterns
This section uses gradient-based analysis to examine the effect of SFT on transformer attention head activation patterns. We identify three key insights: (1) After SFT, LLMs accomplish tasks by selectively activating specific attention heads. (2) Changes in activation patterns for compound tasks can be interpreted as a combination of those in more basic tasks. (3) During SFT, activation patterns exhibit swift changes, even small datasets can cause significant changes.
4.1 Attention Heads are Selectively Activated
@@
In this section, we analyze changes in attention head activation patterns before and after SFT across various tasks and explore the relationships between these changes.
4.1.1 Analytical Method
@@
To analyze changes in model activation patterns before and after SFT for each task, we used three metrics: Gini Coefficient, Coefficient of Variation (CV), and Kurtosis. The Gini Coefficient measures distribution inequality, with values above 0.4 indicating a concentration of activation in a few attention heads. The Coefficient of Variation reflects relative dispersion, with values over 1 signaling substantial variability across attention heads. Kurtosis describes the distribution’s shape, where values above 3 suggest heavy tails and sharp peaks, indicating the presence of extreme values
4.1.2 Experiments Setup
@@
We conducted experiments on models with varying performance levels, including Llama3-8B  AI@Meta (2024), Gemma-7B  Team et al. (2024), and OPT-6.7B  Zhang et al. (2022), across a range of tasks: mathematical reasoning (MATH  Hendrycks et al. (2021b), GSM8K  Cobbe et al. (2021)), coding (Code Search Net  Husain et al. (2019), SGSM  Christ et al. (2024)), and natural language processing and reasoning (HellaSwag  Zellers et al. (2019), Winogrande  Sakaguchi et al. (2021), ARC  Clark et al. (2018)).
4.1.3 Results